# Shapiro Delay Measurements from Fifteen Years of PSR J12311411 Radio Observations

### v1

@@ Abstract
We present 15 years of Nançay and Green Bank radio telescope timing observations for PSR J12311411. This millisecond pulsar is a primary science target for the Neutron Star Interior Composition Explorer telescope (NICER, which discovered its X-ray pulsations), has accumulated near-continuous -ray data since the Fermi-Large Area Telescope’s launch, and has been studied extensively with the Green Bank and Nançay radio telescopes. We have undertaken a campaign with the Green Bank Telescope targeting specific orbital phases designed to improve our constraint on the pulsar’s mass through the detection of a relativistic Shapiro delay. Both frequentist and Bayesian techniques—the latter incorporating priors from white dwarf binary evolution models—are applied to fifteen years of radio observations, yielding relatively weak constraints on the companion and pulsar masses of M⊙ and M⊙, respectively (68.3% CI from Bayesian fits); however, the orbital inclination is measured to better relative precision (°). Restricting the maximum allowed pulsar mass to 3 M⊙ while simultaneously sampling the noise and timing models improves the constraint and lowers the measured mass to M⊙. While our radio-derived inclination result has informed recent NICER X-ray studies of PSR J12311411, the lessons learned from this troublesome pulsar will also bolster future high-precision mass measurement campaigns and resulting constraints on the neutron star interior equation of state.
1 Introduction
@@
Multi-messenger observations in the electromagnetic and gravitational wave bands play a critical role in the effort to determine the equation of state (EoS) of supranuclear-density matter. Observations of relativistic Shapiro delay in radio millisecond pulsars (MSPs) have resulted in precise measurements of the heaviest neutron stars’ masses (e.g. dem10; fon21), imposing a direct constraint on the neutron star interior EoS.
@@
X-ray light curve modeling, facilitated by instruments such as NASA’s Neutron Star Interior Composition Explorer (NICER) telescope (e.g. ril19; mil19; gen17), can constrain the EoS through the simultaneous measurement of masses () and radii (). General relativistic effects from strong gravity affect the X-ray emission from hot spots on the neutron star surface in ways that depend strongly on the stellar compactness. The substantial bending permits viewers to see over the neutron star horizon (bog07). The precise modeling of the observed light curve therefore permits extracting the mass-to-radius () ratio for the neutron star ( and can be individually determined to a certain extent, but an independent measure of improves the estimation of these parameters; an overview is presented in wat16).
@@
One such study by raj21 combines neutron star mass-radius constraints from NICER and the XMM-Newton European Photon Imaging Camera (ril21), radio Shapiro delay observations (cro20; fon21), compact binary mergers GW170817 and GW190425 (abb19a; abb20a), and the kilonova AT2017gfo (kas17). In constraining the mass-to-radius ratio, they find that a 1.4-M⊙ neutron star will have a radius of km or km for piecewise-polytropic and speed-of-sound EoS parameterizations, respectively. Such an analysis is a testament to the power of a multi-messenger observational approach to constraining the neutron star interior EoS.
@@
We present a pulsar timing analysis from radio observations of the MSP J12311411, finding that a significant measurement of Shapiro delay in the radio timing data constrains the pulsar’s mass, albeit to poor precision given the orbital geometry and timing properties of this system. The measurements reported in this manuscript served as conservative priors in NICER’s light curve modeling for this source, which helps break the degeneracy in the M/R ratio and allows discrimination among competing models of hot spot geometry (sal24). Section 2 provides an overview of the Green Bank Telescope (GBT) and Nançay Radio Telescope (NRT) observations, and describes available data from the Fermi-Large Area Telescope (LAT). Section 3 discusses the derivation of pulsar masses via the relativistic Shapiro delay as well as software used in this analysis. Section 4 presents the noise modeling and timing analysis of the data, as well as the various Bayesian approaches taken in deriving the mass of PSR J12311411. A discussion (Section 5) and conclusions (Section 6) follow.
2 Observations & Initial Analysis
@@
J12311411 was discovered with the GBT during a radio pulsar search campaign targeting 25 of the brightest unassociated Fermi-LAT -ray sources111In the Fermi-LAT Bright Source List; see abd09. (ran11). This pulsar is a relatively typical 3.68-ms MSP in a 1.86-day orbit with a white dwarf companion (projected semi-major axis of the pulsar’s orbit = 2.043 lt-s). Since 2011, J12311411 has been intensively observed at radio wavelengths with the NRT, and more sporadically, the GBT (see Figure 1). Although the timing precision of this source is better at 820 MHz than at L-band (1.4 GHz), strong diffractive scintillation sometimes renders this pulsar invisible at lower frequencies. ray19 reported X-ray pulsations from J12311411, with initial spectral modeling suggesting thermal pulsations originating from at least one hot spot on the neutron star’s surface.
2.1 Green Bank Telescope
@@
After its discovery at 820 MHz with the GBT in July 2009, a timing campaign for PSR J12311411 was conducted between December 2009 and August 2010 using the GBT 350-MHz receiver (ran11). Observing epochs were frequent and evenly spaced throughout the entire timing campaign.
@@
We present a new set of observations from the GBT taken in Spring 2019 (project code GBT19A-429) and analyzed in combination with 15 years of additional radio timing data. The 22-hour GBT project comprised five, two-hour observations scheduled at random orbital phases, and two, six-hour observations over superior conjunction (where the Shapiro delay is most prominent). These observations were taken with the 820-MHz receiver and the Green Bank Ultimate Pulsar Processing Instrument backend (GUPPI; dup08) with 200 MHz of bandwidth split into 512 channels, later integrated into 128 channels. All observations were conducted using an online folding mode with coherent dedispersion. Portions of several observations were removed, as diffractive scintillation sometimes rendered the MSP too dim. Notably, the superior conjunction observations saw the MSP scintillated down, reducing the benefit of the orbital-phase-targeted campaign. Calibrated total-intensity data were integrated over 16-minute (1000 s) intervals, and the resulting profiles were matched to a smoothed template using PSRCHIVE (van11) to obtain time-of-arrival (TOA) measurements (see arz18 arz18 for a detailed explanation of this process). The median TOA uncertainty for all GBT data is 6.2 s.
2.2 Nançay Radio Telescope
@@
PSR J12311411 has been observed near-continuously with the NRT since November 2009, with the Berkeley-Orléans-Nançay (BON) backend until August 2011, and with the NUPPI backend after it became the primary pulsar timing instrument in August 2011 (cog13; des11). BON observations were coherently dedispersed and carried out over a bandwidth of 128 MHz. NUPPI observations were also coherently dedispersed and conducted over an increased bandwidth of 512 MHz. The bulk of BON (respectively, NUPPI) observations were conducted at a central frequency of 1398 MHz (1484 MHz). Additional details on the NRT and on pulsar observations with this telescope can be found in gui23 (gui23).
@@
NRT observations were integrated in time, and the resulting profiles were then split into multiple frequency sub-bands. BON observations were split into two sub-bands of 64 MHz each, and NUPPI observations were split into four sub-bands of 128 MHz. Eight observations with the highest signal-to-noise ratios were then smoothed to produce standard profiles for each of the two NRT datasets. The polarization information was preserved throughout these data preparation steps. TOAs were finally extracted from the NRT observations using the Matrix Template Matching (MTM) method of PSRCHIVE (van06), which uses information in all four Stokes parameters to extract TOAs while compensating for differences between the profile and the template.
@@
In addition, as documented in gui23 (gui23), the replacement of a pre-amplifier in 2019 significantly modified the polarimetric response of the NRT, which motivated a revision of the polarization calibration procedure in late 2019. Consequently, all NUPPI observations recorded after mid-2019 were calibrated using the significantly improved calibration procedure, while older NUPPI observations were calibrated with a more generic method. For the purposes of fitting and noise modeling, we therefore divided the NUPPI TOA datasets into two groups: pre- and post-MJD 58600 data, with a JUMP (time-offset parameter) between them. Nevertheless, the MTM method was shown to be effective at reducing timing noise resulting from imperfect polarization calibration (gui23).
2.3 Fermi Large Area Telescope
The Large Area Telescope (LAT) is a GeV -ray instrument aboard NASA’s Fermi satellite, which was launched in June 2008 (atw09). PSR J12311411 is one of the two brightest -ray MSPs as seen from Earth (abd13), and is the best-timed MSP in ker22 ker22, the first data release of the “Gamma Pulsar Timing Array” (GPTA). However, these data are most useful for studying long-timescale phenomena, whereas the radio data sets used in this work are ideal for the study of binary Shapiro delays. For comparison, when -ray data were integrated to produce radio-like TOAs, only two TOAs/year of comparable timing precision (to the thousands of TOAs used in our radio analysis) could be constructed from LAT data. An analysis of Fermi-LAT data for PSR J12311411 will be included in the forthcoming second data release of the GPTA.
3 Methods Overview
Among the many high-precision astrophysical measurements facilitated by pulsar timing is Shapiro delay, a general relativistic effect that is observable in a small subset of highly inclined MSP binaries (sha64). When an MSP passes behind its stellar companion along an observer’s line of sight, the companion’s spacetime deformation causes pulsations from the MSP to arrive later than would be otherwise expected. For circular orbits, the maximum delay occurs at orbital phase = 0.25 (superior conjunction), and is observable as a spike in timing residuals of order 10 s. By measuring the extent and shape of this delay, one obtains the post-Keplerian parameters range () and shape (), the definition of which is dependent on one’s assumed gravitational model. Here, we assume the validity of general relativity; see e.g. tay89 tay89; dam92 dam92. When combined with the Keplerian mass function (Equation 1), measurements of and (Equations 2 and 3, tay92 tay92) uniquely constrain the pulsar and companion masses and :
@@ 3 Methods Overview
This work employs a variety of standard pulsar timing tools to constrain the mass of J12311411. For both Nançay and GBT observations, we use the PINT (luo21) pulsar timing software package. PINT features both traditional fitting routines for pulse times of arrival (TOAs) and Markov Chain Monte Carlo (MCMC) routines. The specifics of these techniques are described in Section 4. Bayesian noise estimation of red and white noise parameters (see Section 4.1) was performed using ENTERPRISE (enterprise). All fits employed the DE440 planetary ephemeris provided by NASA’s Jet Propulsion Laboratory, and used the TDB timescale.
4 Analysis & Results
The data set presented in this work requires careful treatment for a number of reasons. Among these are the aforementioned strong scintillation, heterogeneous composition of backend and receiver combinations, and poor timing precision with significant white (achromatic, stochastic) noise. Covariances also exist between slowly-varying astrometric and general relativistic parameters and red (low-frequency) noise processes, such as DM variations (chromatic red noise) and borderline-significant intrinsic red noise (achromatic red noise; see Section 4.1). Because the Shapiro delay’s timing signature is strongly dependent on the orbital inclination angle, an 80 tilt yields a notably weak Shapiro effect. If PSR J12311411 were instead inclined at = 90, the measurable Shapiro delay would be 3 times larger and far less susceptible to timing and noise model changes.
@@ 4 Analysis & Results
A gridding scheme that performs a least-squares fit with Shapiro delay parameters fixed at each of a set of -sin() pairs and with white and red noise parameter values measured a priori (Section 4.2);
A Bayesian fit with astrophysically-derived priors where all parameters, including the Shapiro delay parameters, are fit simultaneously but noise is measured a priori (Section 4.3);
An exploratory fully-generalized fit where the noise and timing models are determined simultaneously, with some timing parameters being treated as non-linear (Section 4.4).
@@
The “Measured Value” column of Table 2 presents the results of a typical PINT linear regression fit of the combined data set, including the Shapiro delay parameters and sin(). We employed the binary timing model ELL1 (lan01) for this nearly circular MSP system. ELL1 includes the following orbital parameters: projected semi-major axis (); binary orbital period (); epoch of ascending node (TASC); and the two Laplace-Lagrange parameters, = sin () and = cos (), the orbital eccentricity multiplied by the sine and cosine of periastron longitude, respectively. Because these values are already known to high precision, the greatest source of uncertainty in our analyses is measuring the more poorly constrained Shapiro delay parameters and sin(), as well as astrometric (parallax, proper motion) and other post-Keplerian parameters ().
4.1 Noise Analysis for Radio Data
@@
Pulsar timing precision is additionally limited by stochastic white noise sources beyond radiometer noise, such as pulse jitter and radio frequency interference. We contend with these noise sources by including two white noise parameters in our fits for each observing frequency and backend combination included in our analysis (see Table 1). The first is EFAC, a scaling error applied to TOA uncertainties; the second is EQUAD, a term added in quadrature to the TOA uncertainties. More details about the use of these parameters can be found in, e.g., arz18. We note that ECORR, a third commonly used white noise parameter that accounts for inter-channel correlations among TOAs, is not employed here due to a lack of frequency coverage and resolution (lam17).
@@
The white noise parameter estimations were derived using the ENTERPRISE modeling suite, with separate EFAC and EQUAD values measured for each backend-receiver combination. Early GBT 820-MHz observations are treated separately from the 2019 observations, and the NUPPI TOAs are split into two groups based on the significant instrumental changes in 2019. Measured white noise parameter values are presented in Table 1. JUMP parameters are included for all backend-receiver combinations except the pre-58600 subset of NUPPI L-band data, relative to which all other JUMPs are measured. GBT and NRT data have relatively large and small associated offsets, respectively (with the exception of the poorly-constrained JUMP associated with the short 2009 GBT data set). This is expected when choosing the NRT data as the reference.
| Backend & Receiver | EFAC | EQUAD (s) | JUMP () |
| GBT 350 MHz | 1.360 | 0.053 | |
| GBT 820 MHz (2009) | 1.745 | 0.079 | |
| GBT 820 MHz (2019) | 1.146 | 3.838 | |
| NRT L-band, BON (2009) | 0.691 | 0.234 | |
| NRT L-band, NUPPI (pre-58600) | 1.061 | 0.073 | – |
| NRT L-band, NUPPI (post-58600) | 1.064 | 0.124 |
@@
In a subset of well timed MSPs, achromatic red noise (spin noise), likely caused by irregularities in neutron star rotation, can be significantly measured over long time scales. The power spectral density of this red noise is typically modeled with a power law, which has been shown to be sufficient for MSPs (cord16; gon20). A free-spectral model specified over 14 frequency bins was used for exploratory purposes during this work in an effort to pinpoint the origin of relevant noise processes (len16; haz20); however, the results presented here employed a more standard power-law red noise model with 30 frequency bins. Though each run tested for the presence of red noise, none met the inclusion criterion (Bayes factor ¿100) when and parallax were present in the timing model. Section 5 includes a discussion of the impact of red noise model selection on MSP mass measurements.
@@
| a: Values quoted in Section 4.3 should supersede this fit. | |
4.2 PINT Gridding

### v2

@@ Abstract
We present 15 years of Nançay and Green Bank radio telescope timing observations for PSR J12311411. This millisecond pulsar is a primary science target for the Neutron Star Interior Composition Explorer telescope (NICER, which discovered its X-ray pulsations), has accumulated near-continuous -ray data since the Fermi-Large Area Telescope’s launch, and has been studied extensively with the Green Bank and Nançay radio telescopes. We have undertaken a campaign with the Green Bank Telescope targeting specific orbital phases designed to improve our constraint on the pulsar’s mass through the detection of a relativistic Shapiro delay. Both frequentist and Bayesian techniques—the latter incorporating priors from white dwarf binary evolution models—are applied to fifteen years of radio observations, yielding relatively weak constraints on the companion and pulsar masses of M⊙ and M⊙, respectively (68.3% CI from Bayesian fits); however, the orbital inclination is measured to better relative precision (°). Restricting the maximum allowed pulsar mass to 3 M⊙ improves the constraint and lowers the measured mass to M⊙. A fully-generalized Bayesian fit that simultaneously samples the noise and timing models yields a pulsar mass in close agreement with this value. While our radio-derived inclination result has informed recent NICER X-ray studies of PSR J12311411, the lessons learned from this troublesome pulsar will also bolster future high-precision mass measurement campaigns and resulting constraints on the neutron star interior equation of state.
I Introduction
@@
Multi-messenger observations in the electromagnetic and gravitational wave bands play a critical role in the effort to determine the equation of state (EoS) of supranuclear-density matter. Observations of relativistic Shapiro delay in radio millisecond pulsars (MSPs) have resulted in precise measurements of the heaviest neutron stars’ masses (e.g. Demorest et al., 2010; Fonseca et al., 2021), imposing a direct constraint on the neutron star interior EoS.
@@
X-ray light curve modeling, facilitated by instruments such as NASA’s Neutron Star Interior Composition Explorer (NICER) telescope (e.g. Riley et al., 2019; Miller et al., 2019; Gendreau and Arzoumanian, 2017), can constrain the EoS through the simultaneous measurement of masses () and radii (). General relativistic effects from strong gravity affect the X-ray emission from hot spots on the neutron star surface in ways that depend strongly on the stellar compactness. The substantial bending permits viewers to see over the neutron star horizon (Bogdanov et al., 2007). The precise modeling of the observed light curve therefore permits extracting the mass-to-radius () ratio for the neutron star ( and can be individually determined to a certain extent, but an independent measure of improves the estimation of these parameters; an overview is presented in Watts et al., 2016).
@@
One such study by Raaijmakers et al. (2021) combines neutron star mass-radius constraints from NICER and the XMM-Newton European Photon Imaging Camera (Riley et al., 2021), radio Shapiro delay observations (Cromartie et al., 2020; Fonseca et al., 2021), compact binary mergers GW170817 and GW190425 (Abbott et al., 2019, 2020), and the kilonova AT2017gfo (Kasliwal et al., 2017). In constraining the mass-to-radius ratio, they find that a 1.4-M⊙ neutron star will have a radius of km or km for piecewise-polytropic and speed-of-sound EoS parameterizations, respectively. Such an analysis is a testament to the power of a multi-messenger observational approach to constraining the neutron star interior EoS.
@@
We present a pulsar timing analysis from radio observations of the MSP J12311411, finding that a significant measurement of Shapiro delay in the radio timing data constrains the pulsar’s mass, albeit to poor precision given the orbital geometry and timing properties of this system. The measurements reported in this manuscript served as conservative priors in NICER’s light curve modeling for this source, which helps break the degeneracy in the M/R ratio and allows discrimination among competing models of hot spot geometry (Salmi et al., 2024). Section II provides an overview of the Green Bank Telescope (GBT) and Nançay Radio Telescope (NRT) observations, and describes available data from the Fermi-Large Area Telescope (LAT). Section III discusses the derivation of pulsar masses via the relativistic Shapiro delay as well as software used in this analysis. Section IV presents the noise modeling and timing analysis of the data, as well as the various Bayesian approaches taken in deriving the mass of PSR J12311411. A discussion (Section V) and conclusions (Section VI) follow.
II Observations & Initial Analysis
@@
J12311411 was discovered with the GBT during a radio pulsar search campaign targeting 25 of the brightest unassociated Fermi-LAT -ray sources111In the Fermi-LAT Bright Source List; see Abdo et al. (2009). (Ransom et al., 2011). This pulsar is a relatively typical 3.68-ms MSP in a 1.86-day orbit with a white dwarf companion (projected semi-major axis of the pulsar’s orbit = 2.043 lt-s). Since 2011, J12311411 has been intensively observed at radio wavelengths with the NRT, and more sporadically, the GBT (see Figure 1). Although the timing precision of this source is better at 820 MHz than at L-band (1.4 GHz), strong diffractive scintillation sometimes renders this pulsar invisible at lower frequencies. Ray et al. (2019) reported X-ray pulsations from J12311411, with initial spectral modeling suggesting thermal pulsations originating from at least one hot spot on the neutron star’s surface.
II.1 Green Bank Telescope
@@
After its discovery at 820 MHz with the GBT in July 2009, a timing campaign for PSR J12311411 was conducted between December 2009 and August 2010 using the GBT 350-MHz receiver (Ransom et al., 2011). Observing epochs were frequent and evenly spaced throughout the entire timing campaign.
@@
We present a new set of observations from the GBT taken in Spring 2019 (project code GBT19A-429) and analyzed in combination with 15 years of additional radio timing data. The 22-hour GBT project comprised five, two-hour observations scheduled at random orbital phases, and two, six-hour observations over superior conjunction (where the Shapiro delay is most prominent). These observations were taken with the 820-MHz receiver and the Green Bank Ultimate Pulsar Processing Instrument backend (GUPPI; DuPlain et al., 2008) with 200 MHz of bandwidth split into 512 channels, later integrated into 128-channel groups, for a total of 4 sub-bands. All observations were conducted using an online folding mode with coherent dedispersion. Portions of several observations were removed, as diffractive scintillation sometimes rendered the MSP too dim. Notably, the superior conjunction observations saw the MSP scintillated down, reducing the benefit of the orbital-phase-targeted campaign. Calibrated total-intensity data were integrated over 16-minute (1000 s) intervals, and the resulting profiles were matched to a smoothed template using PSRCHIVE (van Straten et al., 2011) to obtain time-of-arrival (TOA) measurements (see Arzoumanian et al. 2018 for a detailed explanation of this process). The median TOA uncertainty for all GBT data is 6.2 s.
II.2 Nançay Radio Telescope
@@
PSR J12311411 has been observed near-continuously with the NRT since November 2009, with the Berkeley-Orléans-Nançay (BON) backend until August 2011, and with the Nançay Ultimate Pulsar Processing Instrument (NUPPI) backend after it became the primary pulsar timing instrument in August 2011 (Cognard et al., 2013; Desvignes et al., 2011). BON observations were coherently dedispersed and carried out over a bandwidth of 128 MHz. NUPPI observations were also coherently dedispersed and conducted over an increased bandwidth of 512 MHz. The bulk of BON (respectively, NUPPI) observations were conducted at a central frequency of 1398 MHz (1484 MHz). Additional details on the NRT and on pulsar observations with this telescope can be found in Guillemot et al. (2023).
@@
NRT observations were integrated in time, and the resulting profiles were then split into multiple frequency sub-bands. BON observations were integrated into two channels of 64 MHz each, and NUPPI observations were integrated into four channels of 128 MHz. Eight observations with the highest signal-to-noise ratios were then smoothed to produce standard profiles for each of the two NRT datasets. The polarization information was preserved throughout these data preparation steps. TOAs were finally extracted from the NRT observations using the Matrix Template Matching (MTM) method of PSRCHIVE (van Straten, 2006), which uses information in all four Stokes parameters to extract TOAs while compensating for differences between the profile and the template.
@@
In addition, as documented in Guillemot et al. (2023), the replacement of a pre-amplifier in 2019 significantly modified the polarimetric response of the NRT, which motivated a revision of the polarization calibration procedure in late 2019. Consequently, all NUPPI observations recorded after mid-2019 were calibrated using the significantly improved calibration procedure, while older NUPPI observations were calibrated with a more generic method. For the purposes of fitting and noise modeling, we therefore divided the NUPPI TOA datasets into two groups: pre- and post-MJD 58600 data, with a JUMP (time-offset parameter) between them. Nevertheless, the MTM method was shown to be effective at reducing timing noise resulting from imperfect polarization calibration (Guillemot et al., 2023).
II.3 Fermi Large Area Telescope
The Large Area Telescope (LAT) is a GeV -ray instrument aboard NASA’s Fermi satellite, which was launched in June 2008 (Atwood et al., 2009). PSR J12311411 is one of the two brightest -ray MSPs as seen from Earth (Abdo et al., 2013), and is the best-timed MSP in FERMI-LAT Collaboration et al. 2022, the first data release of the “Gamma Pulsar Timing Array” (GPTA). However, these data are most useful for studying long-timescale phenomena, whereas the radio data sets used in this work are ideal for the study of binary Shapiro delays. For comparison, when -ray data were integrated to produce radio-like TOAs, only two TOAs/year of comparable timing precision (to the thousands of TOAs used in our radio analysis) could be constructed from LAT data. An analysis of Fermi-LAT data for PSR J12311411 will be included in the forthcoming second data release of the GPTA.
III Methods Overview
Among the many high-precision astrophysical measurements facilitated by pulsar timing is Shapiro delay, a general relativistic effect that is observable in a small subset of highly inclined MSP binaries (Shapiro, 1964). When an MSP passes behind its stellar companion along an observer’s line of sight, the companion’s spacetime deformation causes pulsations from the MSP to arrive later than would be otherwise expected. For circular orbits, the maximum delay occurs at orbital phase = 0.25 (superior conjunction), and is observable as a spike in timing residuals of order 10 s. By measuring the extent and shape of this delay, one obtains the post-Keplerian parameters range () and shape (), the definition of which is dependent on one’s assumed gravitational model. Here, we assume the validity of general relativity; see e.g. Taylor and Weisberg 1989; Damour and Taylor 1992. When combined with the Keplerian mass function (Equation 1), measurements of and (Equations 2 and 3, Taylor et al. 1992) uniquely constrain the pulsar and companion masses and :
@@ III Methods Overview
This work employs a variety of standard pulsar timing tools to constrain the mass of J12311411. For both Nançay and GBT observations, we use the PINT (Luo et al., 2021) pulsar timing software package to fit sub-banded TOAs. PINT features both traditional fitting routines for pulse times of arrival (TOAs) and Markov Chain Monte Carlo (MCMC) routines. The specifics of these techniques are described in Section IV. Bayesian noise estimation of red and white noise parameters (see Section IV.1) was performed using ENTERPRISE (Ellis et al., 2020). All fits employed the DE440 planetary ephemeris provided by NASA’s Jet Propulsion Laboratory, and used the TDB timescale.
IV Analysis & Results
The data set presented in this work requires careful treatment for a number of reasons. Among these are the aforementioned strong scintillation, heterogeneous composition of backend and receiver combinations, and poor timing precision with significant white (achromatic, stochastic) noise. Covariances also exist between slowly-varying astrometric and general relativistic parameters and red (low-frequency) noise processes, such as DM variations (chromatic red noise) and borderline-significant intrinsic red noise (achromatic red noise; see Section IV.1). Because the Shapiro delay’s timing signature is strongly dependent on the orbital inclination angle, an 80 tilt yields a notably weak Shapiro effect. If PSR J12311411 were instead inclined at = 90, the measurable Shapiro delay would be 3 times larger and far less susceptible to timing and noise model changes.
@@ IV Analysis & Results
A gridding scheme that performs a least-squares fit with Shapiro delay parameters fixed at each of a set of -sin() pairs and with white and red noise parameter values measured a priori (Section IV.2);
A Bayesian fit with astrophysically-derived priors where all parameters, including the Shapiro delay parameters, are fit simultaneously but noise is measured a priori (Section IV.3);
An exploratory fully-generalized fit where the noise and timing models are determined simultaneously, with some timing parameters being treated as non-linear (Section IV.4).
@@
The “Measured Value” column of Table 1 presents the results of a typical PINT linear regression fit of the combined data set, including the Shapiro delay parameters and sin(). We employed the binary timing model ELL1 (Lange et al., 2001) for this nearly circular MSP system. ELL1 includes the following orbital parameters: projected semi-major axis (); binary orbital period (); epoch of ascending node (TASC); and the two Laplace-Lagrange parameters, = sin () and = cos (), the orbital eccentricity multiplied by the sine and cosine of periastron longitude, respectively. Because these values are already known to high precision, the greatest source of uncertainty in our analyses is measuring the more poorly constrained Shapiro delay parameters and sin(), as well as astrometric (parallax, proper motion) and other post-Keplerian parameters ().
IV.1 Noise Analysis for Radio Data
@@
Pulsar timing precision is additionally limited by stochastic white noise sources beyond radiometer noise, such as pulse jitter and radio frequency interference. We contend with these noise sources by including two white noise parameters in our fits for each observing frequency and backend combination included in our analysis (see Table 2). The first is EFAC, a scaling error applied to TOA uncertainties; the second is EQUAD, a term added in quadrature to the TOA uncertainties. More details about the use of these parameters can be found in, e.g., Arzoumanian et al. (2018). We note that ECORR, a third commonly used white noise parameter that accounts for inter-channel correlations among TOAs, is not employed here due to a lack of frequency coverage and resolution (Lam et al., 2016).
@@
The white noise parameter estimations were derived using the ENTERPRISE modeling suite and frequency-resolved TOAs, with separate EFAC and EQUAD values measured for each backend-receiver combination. Early GBT 820-MHz observations are treated separately from the 2019 observations, and the NUPPI TOAs are split into two groups based on the significant instrumental changes in 2019. Measured white noise parameter values are presented in Table 2. JUMP parameters are included for all backend-receiver combinations except the pre-58600 subset of NUPPI L-band data, relative to which all other JUMPs are measured. GBT and NRT data have relatively large and small associated offsets, respectively (with the exception of the poorly-constrained JUMP associated with the short 2009 GBT data set). This is expected when choosing the NRT data as the reference.
@@
In a subset of well timed MSPs, achromatic red noise (spin noise), likely caused by irregularities in neutron star rotation, can be significantly measured over long time scales. The power spectral density of this red noise is typically modeled with a power law, which has been shown to be sufficient for MSPs (Cordes et al., 2016; Goncharov et al., 2020). A free-spectral model specified over 14 frequency bins was used for exploratory purposes during this work in an effort to pinpoint the origin of relevant noise processes (Lentati et al., 2016; Hazboun et al., 2020); however, the results presented here employed a more standard power-law red noise model with 30 frequency bins. Though each run tested for the presence of red noise, none met the inclusion criterion (Bayes factor ¿100) when and parallax were present in the timing model. Section V includes a discussion of the impact of red noise model selection on MSP mass measurements.
@@
| a: Values quoted in Section IV.3 should supersede this fit. | |
IV.2 PINT Gridding